News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 18 July 2003 7.00am EDT

During the last week Chandra completed the observing schedule as planned.

The first Aspect Camera dark current calibration following the lowering of the CCD temperature from -10 to -15 deg C was conducted on July 14. An Initial analysis of the data indicates a reduction in the number of warm pixels by 50-60%, as expected.

During the last week the IRU-1 gyro-1 current remained near a base level of 150 mA with some noisy periods. Products and procedures are on track for the planned swap to IRU-2 scheduled for 7.30am EDT on July 19. The swap is planned during a perigee passage and is not expected to impact science operations.

The observing schedule for next week is shown below and includes an observation of Eta Carina coordinated with RXTE.

------------------------------------------
Eta Carinae            ACIS-S/HETG  Jul 19
SCO X-1                ACIS-S/HETG  Jul 21
PKS 2316-423           ACIS-S
PKS0537-28_2           ACIS-I
Radiation Belts
NGC 7212               ACIS-S       Jul 22
NGC 4051               HRC-S/LETG
HZ43                   HRC-S/LETG   Jul 23
HZ43                   HRC-I/LETG   Jul 24
M8                     ACIS-I
Radiation Belts
M8                     ACIS-I       Jul 25

------------------------------------------

All spacecraft subsystems, with the exception of the MUPS thruster A-1 under investigation, continued to operate nominally.

Return to Status Reports